OpenStudy (anonymous):

Can someone help me with substitution?? x^2+4y^2=4 y=x+1 Use the substitution method to solve for the two values of x. I got (0,1) and (-8/5, -3/5). But I feel like that isn't correct.

OpenStudy (anonymous):

first, you should isolate the x in the first equation

OpenStudy (anonymous):

so, x^2= 4y^2-4?

OpenStudy (anonymous):

it has to be in slope intercept form

OpenStudy (anonymous):

first one is an ellipse, second is a line. how many points of intersection are possible?

OpenStudy (anonymous):

two?

OpenStudy (anonymous):

0, 1, 2

OpenStudy (anonymous):

substitute x+1 for y in the first equation and solve for x.

OpenStudy (anonymous):

x^2 + 4(x+1)^2 = 4

OpenStudy (anonymous):

Got it! I'm solving it. Am I correct so far? x^2+4*(x+1)^2=4 x^2+4x^2+16=4 -16 -16 x^2+4x^2= -12?

OpenStudy (anonymous):

no... (x+1)(x+1) = x^2 + 2x + 1 4(x+1)^2 = 4(x^2 + 2x + 1)

OpenStudy (anonymous):

Oh goodness! I'm so lost! D;

OpenStudy (anonymous):

x^2 + 4(x+1)^2 = x^2 + 4(x^2 + 2x + 1) = x^2 + 4x^2 + 8x + 4 = 4 so 5x^2 + 8x = 0 => x(5x + 8) = 0 => x = 0, x = -8/5 those are the x values of the points of intersection. plug those into y = x+1 to find the y values.

OpenStudy (anonymous):

So y=1 and y= -0.6?